# Nonhermitian topological zero modes at smooth domain walls: Exact solutions

### v1

@@ IV.4 Local properties and local topological invariant
In all cases, being featureless and nonfeatureless with short or long hair, the long-distance behavior $x\gg w$ of the zero modes are completely described by Eq. 8 with the exponents $\mu$ and $\kappa$. In particular, the different asymptotic regimes are completely determined by the quantities ${M}_{L,R}$, ${K}_{L,R}$, and $W_{L,R}$ defined in Eqs. 11 and 11c which are in turn determined by the values of the fields ${m}(x)$ and ${v}(x)$ at long distances $x\to\pm\infty$. However, to gain a physical intuition on the wavefunction of the zero modes at short distances $x\lesssim w$, one can introduce the following functions
@@
which generalize the quantities ${M}_{L,R}$, ${K}_{L,R}$, and $W_{L,R}$ to the whole interval $-\infty<x<\infty$. We will refer to $W(x)$ as the local topological invariant and to $M(x)$ as the local topological mass.
In agreement with the bulk boundary correspondence generalization to the nonhermitian case, zero modes on the interval $-\infty<x<\infty$ appear when the topological invariants on the left and on the right of the domain wall are different $W_{L}\neq W_{R}$ or, equivalently, when the local topological invariant $W(x)$ changes sign an odd number of times (if $|W_{L}-W_{R}|=1$) or an even number of times (if $|W_{L}-W_{R}|=2$) on the interval $-\infty<x<\infty$. The points $x_{i}^{*}$ where the local topological invariant changes must be $|x_{i}^{*}|\lesssim w$, since the fields become uniform at larger distances, whereby the local topological invariant becomes as well. One can infer from the bulk-boundary correspondence and its nonhermitian generalization that, in this case, the zero modes will be localized in correspondence with these points $x\approx x_{i}^{*}$. Analogously, zero modes on the interval $0\leq x<\infty$ appear when the topological invariant on the right $W_{R}=W(x\to\infty)$ is nonzero. If the local topological invariant $W(x)$ is constant, then the zero modes will localize at $x=0$. However, if the local topological invariant $W(x)$ changes at the points $x_{i}^{*}$, then the zero modes will localize at the points $x\approx x_{i}^{*}$.
On the other hand, the quantity ${K}_{L,R}$ determines whether the zero modes shows exponentially damped oscillations (${K}_{L,R}\neq 0$) or a simple exponential decay without oscillations (${K}_{L,R}=0$) at large distances Similarly, the value of the function ${K}(x)$ will indicate the presence of oscillations for ${K}(x)\neq 0$ and the absence of oscillations for ${K}(x)=0$ at any given point $x$.
@@ IV.5 Examples
In this Section, we show some examples of zero modes in the nonhermitian case in Eqs. 7 and 9. Figure 2 show zero modes on the interval $-\infty<x<\infty$ in the case where the width of the smooth domain wall $w$ is smaller than or comparable with the characteristic decay lengths $\xi$, for different choices of the parameters ${m}_{0,1,2}$ and ${v}_{1,2}$. The exponential decay is visible with the exception of a small region for $|x|<w<\xi$: These modes have "short hair", i.e., they are nonfeatureless for $|x|\lesssim w$ but can be regarded as featureless for distances $|x|\gtrsim w$. Figure 3 shows analogous cases, but where the width of the smooth domain wall $w$ is larger than the characteristic decay lengths $\xi$. The most noticeable difference is that the exponential decay of the modes is now visible only at large distances $|x|\gtrsim w>\xi$: These modes have "long hair", i.e., they are nonfeatureless at finite length scales. Figures 4 and 5 shows zero modes on the interval $0\leq x<\infty$.
The local features of the zero modes become prominent when $w\gg\xi$ (long hair) and when the field around the domain wall is large. More remarkably, the zero modes on Fig. 3(c) and Fig. 3(d) and on Fig. 5(d) are localized far away from the center of the domain wall in correspondence of a point $|x_{1}^{*}|>\xi$ where the local topological invariant $W(x)$ changes and the local topological mass ${M}(x)$ changes its sign.
@@
Figure 2: Zero modes on the interval $-\infty<x<\infty$ with a smooth domain wall given by S-shaped, symmetric or asymmetric Pöschl–Teller, or constant fields ${m}(x)$ and ${v}(x)$, and with its width smaller than or comparable with the characteristic decay lengths $w\lesssim\xi$. Panels on different rows from top to bottom show the spatial dependence of i) the real and imaginary parts of the wavefunctions, ii) the natural logarithm of the norm of the wavefunctions, iii) the real and imaginary parts of the fields ${m}(x)$ and ${v}(x)$, iv) the quantities ${M}(x)$ and ${K}(x)$. Panels on different columns from left to right correspond to different choices of the parameters ${m}_{0,1,2}$ and ${v}_{1,2}$, giving (a) single mode with exponentially damped oscillations on the right side and exponential decay on the left side of a smooth domain wall with an S-shaped field ${m}(x)$ and a complex S-shaped field ${v}(x)$ with $\mathrm{Im}({v}_{L})\neq 0$ (nonhermitian), (b) single mode with exponentially damped oscillations and nonuniform complex phase on the right side and exponential decay on the left side of a smooth domain wall with complex S-shaped field ${m}(x)$ with a nonzero imaginary component (nonhermitian). (c) single mode with exponentially damped oscillations and nonuniform complex phase on the right side and exponential decay on the left side of a smooth domain wall with asymmetric Pöschl–Teller field ${m}(x)$ and a complex S-shaped field ${v}(x)$ with $\mathrm{Im}({v}_{L})\neq 0$ (nonhermitian), (d) single mode with exponentially damped oscillations and nonuniform complex phase on the right side and exponential decay on the left side of a smooth domain wall with complex asymmetric Pöschl–Teller field ${m}(x)$ with a nonzero imaginary component (nonhermitian).
Figure 3: Zero modes on the interval $-\infty<x<\infty$ with S-shaped, symmetric, or asymmetric Pöschl–Teller, or constant fields ${m}(x)$ and ${v}(x)$ as in Fig. 2, but where the width of the smooth domain wall is larger than the characteristic decay lengths $w>\xi$.
Figure 4: Zero modes on the interval $0\leq x<\infty$ with a smooth domain wall given by S-shaped, symmetric or asymmetric Pöschl–Teller, or constant fields ${m}(x)$ and ${v}(x)$, and with $w\lesssim\xi$. Panels on different rows are as in Fig. 2. Panels on different columns from left to right correspond to different choices of the parameters ${m}_{0,1,2}$ and ${v}_{1,2}$ analogous to Fig. 2, giving (a) single mode with exponentially damped oscillations with nonuniform complex phase on the right side of a smooth domain wall with S-shaped fields ${m}(x)$ and ${v}(x)$ with $\mathrm{Im}({v}_{R})\neq 0$ (nonhermitian), (b) single mode with exponentially damped oscillations with nonuniform complex phase on the right side of a smooth domain wall with S-shaped field ${m}(x)$ with a nonzero imaginary component (nonhermitian). (c) single mode with exponentially damped oscillations with nonuniform complex phase on the right side of a smooth domain wall with asymmetric Pöschl–Teller field ${m}(x)$ and S-shaped field ${v}(x)$ with $\mathrm{Im}({v}_{R})\neq 0$ (nonhermitian), (d) single mode with exponentially damped oscillations with nonuniform complex phase on the right side of a smooth domain wall with asymmetric Pöschl–Teller field ${m}(x)$ with a nonzero imaginary component (nonhermitian).
@@
Figure 5: Zero modes on the interval $0\leq x<\infty$ with S-shaped, symmetric or asymmetric Pöschl–Teller, or constant fields ${m}(x)$ and ${v}(x)$ as in Fig. 4, but with $w>\xi$.
@@ IV.6 Experimental detection
which is valid for both hermitian and nonhermitian cases and both featureless and nonfeatureless zero modes. In the hermitian or in the asymptotically hermitian cases (i.e., if ${v}_{L,R},{m}_{L,R}\in\mathbb{R}$) and if the modes exhibit exponentially damped oscillations $\kappa_{L,R}\neq 0$, one gets $\mu_{L,R}=\pm{v}_{L,R}$ and consequently
@@
which is valid for both featureless and nonfeatureless zero modes. Note that all the quantities on the left hand of Eqs. 30 and 31 describe the wavefunction at large distances, in principle measurable by spatially-resolved spectroscopies, while the quantities on the right hand depend on the "bulk" properties of the system, i.e., the dispersion of bulk excitations of the system at large distances $x/w\to\pm\infty$. Thus, one can design an experiment in which some external parameters are varied and simultaneously measure the bulk dispersion at large distances and the decay rate and momenta of the zero modes. From these two sets of independent measures, one can thus obtain the quantities ${m}_{L,R}^{\text{exp}}$, ${v}_{L,R}^{\text{exp}}$ (bulk excitations) and $\mu_{L,R}^{\text{exp}}$, $\kappa_{L,R}^{\text{exp}}$ (zero modes). Strong evidence of the presence of zero modes is obtained when these quantities satisfy Eq. 30 on an extended range of external parameters.
@@ IV.7 Topological phases and differential order
We note that a general statement can be made on the maximum number of topologically protected modes and the maximum number of topologically distinct phases only by looking at the differential order of the Hamiltonian. The order of the differential equation determined by the Hamiltonian coincides with the maximum number of possible topologically protected modes and with twice the maximum possible value of the topological invariant as a consequence of the bulk-boundary correspondence. A second-order differential equation can support at most 2 linearly independent solutions satisfying the boundary conditions. Therefore, any Hamiltonian that can be reduced to a scalar second-order differential equation as in the modified Jackiw-Rebbi equation in Eq. 3 may exhibit at most 2 topologically protected edge modes. In this case, the critical lines between topologically distinct phases, where the edge modes localize, can only separate phases with topological invariants $W,W^{\prime}$ with $|W-W^{\prime}|\leq 2$, due to the bulk-boundary correspondence. Hence, the topological invariant must be $|W|\leq 1$, with a maximum number of 2 nontrivial phases, regardless of the presence of antiunitary and unitary symmetries. More generally, given that an $n$-order differential equation can support at most $n$ linearly independent solutions satisfying the boundary conditions, any Hamiltonian that can be reduced to a scalar $n$-order differential equation may exhibit at most $n$ topologically-protected edge modes, and the critical lines can only separate distinct phases with topological invariants $W,W^{\prime}$ with $|W-W^{\prime}|\leq n$, which allows the presence of at most $\lfloor n/2\rfloor$ topologically nontrivial phases with $2|W|\leq n$, regardless of the presence of antiunitary and unitary symmetries.
@@ V Conclusions
Concluding, we derived the analytical form of the wavefunctions of nonhermitian topological zero modes localized at smooth domain walls, by extending the modified Jackiw-Rebbi equation to nonhermitian systems with line gaps. Using these analytical results, we showed that the nonhermitian bulk-boundary correspondence for these systems can be directly derived by analyzing the asymptotic and localization properties of the zero modes. We thus identified a universal relation between the decay rates and oscillation wavelengths of the nonhermitian topological zero modes and the underlying scalar fields. This relation provides a direct link between the bulk properties and experimentally measurable physical quantities, which are, in principle, accessible via spatially resolved spectroscopies, suggesting an experimental way to probe nonhermitian topological zero modes. Analogously to their hermitian counterpart, the nonhermitian topological zero modes exhibit different localization behaviors, from featureless zero modes in the sharp domain wall limit to modes with "short" or "long hair" in the presence of smooth domain walls. Additionally, we showed that these modes can exhibit either pure exponential decay or exponentially damped oscillations in different regimes. These findings allow us to understand the physical properties of the zero modes, such as localization and oscillating behavior, that cannot be understood by topological considerations alone.
@@ Acknowledgements.
P. M. is partially supported by the Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Early-Career Scientists KAKENHI Grant No. 23K13028 and No. 20K14375, Grant-in-Aid for Transformative Research Areas (A) KAKENHI Grant No. 22H05111, and Grant-in-Aid for Transformative Research Areas (B) KAKENHI Grant No. 24H00826.
@@ Appendix A Formal solutions of the modified Jackiw-Rebbi equation
We are specifically interested in the cases where the fields ${m}(x)$ and ${v}(x)$ are approximately uniform for $|x|>w$, and present nonnegligible spatial variations for $|x|<w$. This setup describes the case of a smooth domain wall between two uniform phases at $x\ll-w$ and $x\gg w$, or a sharp domain wall between two uniform phases at $x<0$ and $x>0$ in the limit $w\to 0$. In this setup, most of the interesting physics behavior is concentrated near the domain wall, and not at large distances. In this spirit, we map the whole real line $-\infty<x<\infty$ into the finite segment $10<y<$ 1 with the substitution
@@
Here, the characteristic length $w$ is, in principle, arbitrary but will be chosen, in practice, to match the characteristic length scale of the spatial variations of ${m}(x)$ and ${v}(x)$. The modified Jackiw-Rebbi equation in Eq. 3 becomes
@@
are functions of $y$. There are four different choices of the exponents $(\alpha_{L}^{\pm},\alpha_{R}^{\pm})$, and for any of these choices, we obtain a distinct second-order differential equation for the function $F(y)$, which therefore admits two linearly independent solutions $F_{1,2}$. However, since the modified Jackiw-Rebbi is a differential equation is of second order, one can choose at most two solutions that are linearly independent: The general solution of the equation in Eq. 3 is thus given by a linear combination of these two linearly independent solutions, e.g.,
@@
where $F_{1}(y)$ is one of the two solutions of Eq. 38 corresponding to the exponents $(\alpha_{L}^{+},\alpha_{R}^{-})$ and $F_{2}(y)$ is one of the two solutions corresponding to the exponents $(\alpha_{L}^{-},\alpha_{R}^{+})$, chosen such that $\varphi_{1,2}^{s}(y)$ are linearly independent and well behaved at $y=0,1$, i.e., $\varphi_{1,2}^{s}(y\to 0)\sim y^{w\alpha_{L}^{\pm}}$ and $\varphi_{1,2}^{s}(y\to 1)\sim(1-y)^{w\alpha_{R}^{\mp}}$ in order to satisfy the boundary conditions. The last condition mandates that the functions converge in absolute value $|F(y)|<\infty$ for $y=0,1$ or, in the worst case, diverge logarithmically for $y\to 0$ or $y\to 1$. The solutions $F_{1,2}$ of Eq. 38 can be determined by expanding the function $F(y)$ and the fields ${m}(y)$ and ${v}(y)$ in powers of $y$. We note again that the choice of the two linearly independent solutions in Eq. 40 is not unique. Furthermore, choosing two linearly independent solutions is possible even if the exponents are degenerate, i.e., when $\alpha_{L,R}^{+}=\alpha_{L,R}^{-}$, as long as the functions $F_{1,2}$ are linearly independent.
Transforming back to the real line $-\infty<x<\infty$ the general solution is any linear combination
@@
up to normalization constants with exponents given by Eq. 37 and pseudospin in Eq. 2 determined by the boundary conditions, where we define $G({\alpha_{L}},{\alpha_{R}},x)=y(x)^{w{\alpha_{L}}}(1-y(x))^{w{\alpha_{R}}}$ which is thus given by
@@
The asymptotic behavior of the functions $G({\alpha_{L}},{\alpha_{R}},x\to\pm\infty)$ does not depend on $w$ and on the functional form of the fields ${m}(x)$ and ${v}(x)$, but only on their values ${m}_{L,R}$ and ${v}_{L,R}$ at $x\to\pm\infty$. We will show that this property also holds for the solutions $\varphi_{1,2}^{s}(x)$ of the modified Jackiw-Rebbi equation in Eq. 3.

### v2

@@ IV.4 Local properties and local topological invariant
In both featureless and nonfeatureless cases, the long-distance behavior $|x|\gg w$ of the zero modes is completely described by Eq.˜7. In particular, the different asymptotic regimes are completely determined by the quantities ${M}_{L,R}$, ${K}_{L,R}$, and $W_{L,R}$ defined in Eq.˜10 which are in turn determined by the asymtpotic values of the fields m(x){m}(x) and v(x){v}(x) at long distances $x\to\pm\infty$. However, to understand the properties of the wavefunction at short distances $|x|\lesssim w$, we introduce the following functions
@@
which generalize the quantities ${M}_{L,R}$, ${K}_{L,R}$, and $W_{L,R}$. We will refer to $W(x)$ as the local topological invariant and to $M(x)$ as the local topological mass.
In agreement with the bulk boundary correspondence generalization to the nonhermitian case, zero modes on $(-\infty,\infty)$ appear when the left and right topological invariants differ $W_{L}\neq W_{R}$ or, equivalently, when the local topological invariant $W(x)$ changes sign an odd number of times (if $|W_{L}-W_{R}|=1$) or an even number of times (if $|W_{L}-W_{R}|=2$) on the interval $(-\infty,\infty)$. The points $x_{i}^{*}$ where the local topological invariant changes must be $|x_{i}^{*}|\lesssim w$, since the fields become uniform at larger distances, whereby the local topological invariant becomes uniform as well. One can infer from the bulk-boundary correspondence and its nonhermitian generalization that, in this case, the zero modes will be localized in close proximity to these points $x\approx x_{i}^{*}$. Analogously, zero modes on $0,\infty)$ appear when the topological invariant on the right $W_{R}=W(x\to\infty)$ is nonzero. If the local topological invariant $W(x)$ is constant, then the zero modes will localize at $x=0$. However, if the local topological invariant $W(x)$ changes at the points $x_{i}^{*}$, then the zero modes will localize in proximity to the points $x\approx x_{i}^{*}$.
On the other hand, the quantity ${K}_{L,R}$ distinguishes between exponentially damped oscillations (${K}_{L,R}\neq 0$) and exponential decay without oscillations (${K}_{L,R}=0$) at large distances. Similarly, the value of the function K(x){K}(x) will indicate the presence of oscillations for ${K}(x)\neq 0$ and the absence of oscillations for K(x)=0{K}(x)=0 at any given point $x$.
@@ IV.5 Examples
In this Section, we show some examples of zero modes in the nonhermitian case in Eqs.˜6 and 8. Figure˜2 shows nonfeatureless zero modes with "short hair" on $(-\infty,\infty)$ with decay lengths $\xi$ larger than the smooth domain wall width $w$. Figure˜3 shows instead nonfeatureless zero modes with "long hair" with decay lengths $\xi$ smaller than the smooth domain wall width $w$. Figures˜4 and 5 shows zero modes on the interval $0,\infty)$. The local features are more visible in the long hair regime. For instance, the zero modes on Fig.˜3(c) and Fig.˜3(d) and on Fig.˜5(d) do not localize within the smooth the domain wall but in proximity to a point $|x_{1}^{*}|>\xi$ where the local topological invariant $W(x)$ and the local topological mass M(x){M}(x) change.
@@
Figure 2: Zero modes with short hair ($w<\xi$) on $(-\infty,\infty)$ corresponding to a smooth domain wall given by S-shaped, symmetric or asymmetric Pöschl–Teller, or constant fields m(x){m}(x) and v(x){v}(x). Panels on different rows from top to bottom show the spatial dependence of i) the real and imaginary parts of the wavefunctions, ii) the natural logarithm of the norm of the wavefunctions, iii) the real and imaginary parts of the fields m(x){m}(x) and v(x){v}(x), iv) the quantities M(x){M}(x) and K(x){K}(x). Panels on different columns from left to right correspond to different values of ${m}_{0,1,2}$ and ${v}_{1,2}$, giving (a) single mode with exponentially damped oscillations on the right and exponential decay on the left for an S-shaped field m(x){m}(x) and a complex S-shaped field v(x){v}(x) with $\mathrm{Im}({v}_{L})\neq 0$ (nonhermitian), (b) single mode with exponentially damped oscillations and nonuniform complex phase on the right and exponential decay on the left for a complex S-shaped field m(x){m}(x) with a nonzero imaginary component (nonhermitian). (c) single mode with exponentially damped oscillations and nonuniform complex phase on the right and exponential decay on the left for an asymmetric Pöschl–Teller field m(x){m}(x) and a complex S-shaped field v(x){v}(x) with $\mathrm{Im}({v}_{L})\neq 0$ (nonhermitian), (d) single mode with exponentially damped oscillations and nonuniform complex phase on the right and exponential decay on the left for a complex asymmetric Pöschl–Teller field m(x){m}(x) with a nonzero imaginary component (nonhermitian).
@@
Figure 3: Zero modes with long hair ($w>\xi$) on the interval $(-\infty,\infty)$ with S-shaped, symmetric, or asymmetric Pöschl–Teller, or constant fields m(x){m}(x) and v(x){v}(x) as in Fig.˜2.
Figure 4: Zero modes with short hair ($w<\xi$) on the interval $0,\infty)$ with a smooth domain wall given by S-shaped, symmetric or asymmetric Pöschl–Teller, or constant fields m(x){m}(x) and v(x){v}(x),. Panels on different rows are as in Fig.˜2. Panels on different columns from left to right correspond to different values of ${m}_{0,1,2}$ and ${v}_{1,2}$ analogous to Fig.˜2, giving (a) single mode with exponentially damped oscillations with nonuniform complex phase for an S-shaped fields m(x){m}(x) and v(x){v}(x) with $\mathrm{Im}({v}_{R})\neq 0$ (nonhermitian), (b) single mode with exponentially damped oscillations with nonuniform complex phase for an S-shaped field m(x){m}(x) with a nonzero imaginary component (nonhermitian). (c) single mode with exponentially damped oscillations with nonuniform complex phase for an asymmetric Pöschl–Teller field m(x){m}(x) and S-shaped field v(x){v}(x) with $\mathrm{Im}({v}_{R})\neq 0$ (nonhermitian), (d) single mode with exponentially damped oscillations with nonuniform complex phase for an asymmetric Pöschl–Teller field m(x){m}(x) with a nonzero imaginary component (nonhermitian).
Figure 5: Zero modes with long hair ($w>\xi$) on the interval $0,\infty)$ with S-shaped, symmetric or asymmetric Pöschl–Teller, or constant fields m(x){m}(x) and v(x){v}(x) as in Fig.˜4.
@@ IV.6 Experimental detection
which is valid for both hermitian and nonhermitian cases, and both featureless and nonfeatureless zero modes. In the hermitian or in the asymptotically hermitian cases (i.e., if ${v}_{L,R},{m}_{L,R}\in\mathbb{R}$) and in the regime of exponentially damped oscillations $\kappa_{L,R}\neq 0$, one gets $\mu_{L,R}=\pm{v}_{L,R}$ and consequently
@@
for both featureless and nonfeatureless modes. Note that the left-hand side of Eqs.˜30 and 31 contains quantities which characterize the wavefunction at large distances, while the right-hand side contains quantities which depend on the "bulk" properties of the system. Since all these quantities are measurable, one can design an experimental protocol where the bulk dispersion at large distances and the decay rate and momenta of the zero modes are measured independently. We stress the fact that this experimental protocol requires measuring the probability density (and wavefunction) of the eigenmodes and not their mere presence, as in conventional tunneling spectroscopies. Strong evidence of zero modes is obtained when these quantities satisfy Eq.˜30 on an extended range of external parameters.
@@ IV.7 Topological phases and differential order
We note that a general statement can be made on the maximum number of topologically protected modes and the maximum number of topologically distinct phases only by looking at the differential order of the Hamiltonian. The order of the differential equation determined by the Hamiltonian coincides with the maximum number of possible topologically protected modes and with twice the maximum possible value of the topological invariant as a consequence of the bulk-boundary correspondence. A second-order differential equation can support at most 2 linearly independent solutions satisfying the boundary conditions. Therefore, any Hamiltonian that can be reduced to a scalar second-order differential equation as in the modified Jackiw-Rebbi equation in Eq.˜2 may exhibit at most 2 topologically protected edge modes. In this case, the critical lines between topologically distinct phases, where the edge modes localize, can only separate phases with topological invariants $W,W^{\prime}$ with $|W-W^{\prime}|\leq 2$, due to the bulk-boundary correspondence. Hence, the topological invariant must be $|W|\leq 1$, with a maximum number of 2 nontrivial phases, regardless of the presence of antiunitary and unitary symmetries. More generally, given that an $n$-order differential equation can support at most $n$ linearly independent solutions satisfying the boundary conditions, any Hamiltonian that can be reduced to a scalar $n$-order differential equation may exhibit at most $n$ topologically-protected edge modes, and the critical lines can only separate distinct phases with topological invariants $W,W^{\prime}$ with $|W-W^{\prime}|\leq n$, which allows the presence of at most $\lfloor n/2\rfloor$ topologically nontrivial phases with $2|W|\leq n$, regardless of the presence of antiunitary and unitary symmetries.
@@ V Conclusions
Concluding, we derived the analytical form of the wavefunctions of nonhermitian topological zero modes localized at smooth domain walls, by extending the modified Jackiw-Rebbi equation to nonhermitian systems with line gaps. Using these analytical results, we showed that the nonhermitian bulk-boundary correspondence for these systems can be directly derived by analyzing the asymptotic and localization properties of the zero modes. We thus identified a universal relation between the decay rates and oscillation wavelengths of the nonhermitian topological zero modes and the underlying scalar fields. This relation provides a direct link between the bulk properties and experimentally measurable physical quantities, which are, in principle, accessible via spatially resolved spectroscopies, suggesting an experimental way to probe nonhermitian topological zero modes. Analogously to their hermitian counterpart, the nonhermitian topological zero modes exhibit different localization behaviors, from featureless zero modes in the sharp domain wall limit to modes with "short" or "long hair" in the presence of smooth domain walls. Additionally, we showed that these modes can exhibit either pure exponential decay or exponentially damped oscillations in different regimes. These findings allow us to understand the physical properties of the zero modes, such as localization and oscillating behavior, that cannot be understood by topological considerations alone. Notice also that our derivation of the existence of zero-energy modes in the nonhermitian regime (and in the hermitian regime as a special case) does not rely on topological considerations, but on the analysis of the asymptotic behavior of the solutions and, in particular, on their decay rates. Hence, our approach is inherently more general since it also applies to regimes where bulk topological invariants are ill-defined, such as in the case where the rate of variation of the fields is comparable to the localization length of the modes. Indeed, previous works on the bulk-boundary correspondence in nonhermitian systems dealt with energy modes localized at sharp domain walls between topologically inequivalent phases shen_topological_2018 ; gong_topological_2018 ; kawabata_symmetry_2019 . Here, we go beyond this simplifying assumption and provide an analytical and formal approach to deal with the more general and realistic case of smooth domain walls. Obviously, the sharp domain wall regime can be obtained as a special case of our analysis. In this sense, our work generalizes the bulk-boundary correspondence to smooth domain walls.
Finally, we remark that these findings apply to a large range of systems, including topological insulators and superconductors in the hermitian or nonhermitian regimes, and in particular superconductors with $p$-wave pairing kitaev_unpaired_2001 , but can be, in principle, extended to superconductors with $s$-wave pairing oreg_helical_2010 ; lutchyn_majorana_2010 , $d$-wave pairing tanaka_theory_1995 ; kashiwaya_tunnelling_2000 , or odd-frequency pairing tanaka_symmetry_2012 , and may have relevance for the study of the Dirac equation with domain walls in high-energy physics stojkovic_fermionic_2000 .
@@ Acknowledgements.
The authors thank Masatoshi Sato and Yukio Takaka for their useful suggestions and feedback. P. M. is partially supported by the Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Early-Career Scientists KAKENHI Grant No. 23K13028 and No. 20K14375, Grant-in-Aid for Transformative Research Areas (B) KAKENHI Grant No. 24H00826, and Grant-in-Aid for Transformative Research Areas (A) KAKENHI Grant No. 22H05111.
@@ Appendix A Formal solutions of the modified Jackiw-Rebbi equation
We are specifically interested in the cases where the fields m(x){m}(x) and v(x){v}(x) are approximately uniform for $|x|>w$, and present nonnegligible spatial variations for $|x|<w$. This setup describes the case of a smooth domain wall between two uniform phases at $x\ll-w$ and $x\gg w$, or a sharp domain wall between two uniform phases at $x<0$ and $x>0$ in the limit $w\to 0$. In this setup, most of the interesting physics behavior is concentrated near the domain wall, and not at large distances. In this spirit, we map the whole real line $(-\infty,\infty)$ into the finite segment $(0,1)$ with the substitution
@@
Here, the characteristic length $w$ is, in principle, arbitrary but will be chosen, in practice, to match the characteristic length scale of the spatial variations of m(x){m}(x) and v(x){v}(x). The modified Jackiw-Rebbi equation in Eq.˜2 becomes
@@
are functions of $y$. There are four different choices of the exponents $(\alpha_{L}^{\pm},\alpha_{R}^{\pm})$, and for any of these choices, we obtain a distinct second-order differential equation for the function $F(y)$, which therefore admits two linearly independent solutions $F_{1,2}$. However, since the modified Jackiw-Rebbi is a differential equation of second order, one can choose at most two solutions that are linearly independent: The general solution of the equation in Eq.˜2 is thus given by a linear combination of these two linearly independent solutions, e.g.,
@@
where $F_{1}(y)$ is one of the two solutions of Eq.˜38 corresponding to the exponents $(\alpha_{L}^{+},\alpha_{R}^{-})$ and $F_{2}(y)$ is one of the two solutions corresponding to the exponents $(\alpha_{L}^{-},\alpha_{R}^{+})$, chosen such that $\varphi_{1,2}^{s}(y)$ are linearly independent and well behaved at $y=0,1$, i.e., $\varphi_{1,2}^{s}(y\to 0)\sim y^{w\alpha_{L}^{\pm}}$ and $\varphi_{1,2}^{s}(y\to 1)\sim(1-y)^{w\alpha_{R}^{\mp}}$ in order to satisfy the boundary conditions. The last condition mandates that the functions converge in absolute value $|F(y)|<\infty$ for $y=0,1$ or, in the worst case, diverge logarithmically for $y\to 0$ or $y\to 1$. The solutions $F_{1,2}$ of Eq.˜38 can be determined by expanding the function $F(y)$ and the fields m(y){m}(y) and v(y){v}(y) in powers of $y$. We note again that the choice of the two linearly independent solutions in Eq.˜40 is not unique. Furthermore, choosing two linearly independent solutions is possible even if the exponents are degenerate, i.e., when $\alpha_{L,R}^{+}=\alpha_{L,R}^{-}$, as long as the functions $F_{1,2}$ are linearly independent.
Transforming back to the real line $(-\infty,\infty)$ the general solution is any linear combination
@@
up to normalization constants with exponents given by Eq.˜37 and pseudospin determined by the boundary conditions, where we define $G({\alpha_{L}},{\alpha_{R}},x)=y(x)^{w{\alpha_{L}}}(1-y(x))^{w{\alpha_{R}}}$ which is thus given by
@@
The asymptotic behavior of the functions $G({\alpha_{L}},{\alpha_{R}},x\to\pm\infty)$ does not depend on $w$ and on the functional form of the fields m(x){m}(x) and v(x){v}(x), but only on their values ${m}_{L,R}$ and ${v}_{L,R}$ at $x\to\pm\infty$. We will show that this property also holds for the solutions $\varphi_{1,2}^{s}(x)$ of the modified Jackiw-Rebbi equation in Eq.˜2.